OpenStudy (anonymous):

Can someone walk me through this!!!! Alex is ordering salads and drinks for his friends. Salads cost $8 each, drinks cost $4 each, and there is a $6 delivery charge per order. He has $70. If he buys 4 salads, how many drinks can he buy? Solve the inequality 8x + 4y + 6<=70, where x = number of salads and y = number of drinks, to answer the question.

OpenStudy (jdoe0001):

so... say if we had instead " \(\bf 8x + 4y + 6=70 \) " what would it give you if you solve it for "y" ?

OpenStudy (anonymous):

dont we need somethimg for x first?

OpenStudy (anonymous):

8x+4y=64

OpenStudy (adilalvi):

8x + 4y + 6<=70 8x+4y+6=70 8x+4y=70-6 8x+4y=64 now we have given that x=4 number of salad and we need to find y=number of drinks so putt x=4 in last equation. 8(4)+4y=64 32+4y=64 4y=64-32 4y=32 y=32/4 y=8 is the number of drinks.